You have a fixed monthly salary and a commission of $6 per item you sell. If your total earnings at the end of a month after selling 200 items is $2,500, what is your fixed monthly salary?
Answer by dkppathak(439)

You have a fixed monthly salary and a commission of $6 per item you sell. If your total earnings at the end of a month after selling 200 items is $2,500, what is your fixed monthly salary?
let fixed salary is $ x
number of item sold =200
total comission earne for 200 items at the rate of $6 =6x200=1200
X+1200=2500
x=2500-1200=1300
answer $1300